PRODUCT BRANCHING RATIOS OF THE REACTION OF CO WITH H_{3}^{+} AND ITS ISOTOPOMERS

\underline{HUA-GEN YU}, Department of Chemistry, Brookhaven National Laboratory, Upton, NY 11973.

The reaction of CO with H_{3}^{+}, H_{2}D^{+} and HD_{2}^{+} has been studied using a direct ab initio molecular dynamics method, where the energies and forces used in trajectory propagations are determined by a SAC (scaling all correlation)-MP2/cc-pVTZ theory. For the H_{3}^{+} + CO reaction, there are two product channels: (H_{2} + HCO^{+}) and (H_{2} + HOC^{+}). At room temperature, the thermal rate coefficient is predicted to be 1.37×10^{−9} cm^{3}.molecule^{−1}.s^{−1} with a product branching ratio [HOC^{+}]/[HCO^{+}]=0.28. In addition, dynamics results for the CO + H_{2}D^{+}/HD_{2}^{+} reactions will also be reported.